# Model-Agnostic, Temperature-Informed Sampling Enhances Cross-Year Crop Mapping with Deep Learning

### v2

@@ 5 Experiment & Results
Figure 9: Low data regime performance showing classification (a) accuracy and (b) Expected Calibration Error averaged over six folds when training on 10% of the data, comparing T3S (Ours) and U-TAE.
Figure 10: In‐season classification performance at the 75th‐ and 50th‐percentile cutoffs, corresponding to the end of September and the end of June, respectively, showing (a) accuracy and (b) Expected Calibration Error, averaged over six folds, for T3S (Ours) versus U-TAE.
@@
Figure 11: Qualitative comparison for the proposed T3S and U-TAE. In performance maps, red indicates a wrong prediction while green represents a correct prediction. High uncertainty is indicated by yellow, while low is indicated with dark blue.
6 Discussion
Our experiments demonstrate that embedding GDD-informed sampling into a deep learning pipeline not only improves crop classification accuracy but also yields better-calibrated uncertainty estimates compared to leading baselines. This advantage grows when training data are scarce, highlighting the value of domain expertise under low-data regimes.
Injecting domain knowledge into a neural network, i.e., inductive bias, constrains the hypothesis space to solutions that ’make sense’ under known physical or biological rules and this prevents the model from assigning high confidence to predictions when evidence is weak or contradictory (Mitchell, (1980), Ahmadi Daryakenari et al., (2024)). In our case inductive bias is introduced via GDD-informed sampling. By incorporating thermal time alignment, we align satellite time series with actual crop phenology. This inductive bias acts as a regularizing prior, steering the model toward meaningful growth stages and away from noisy or cloud-contaminated observations. As a result, high-confidence predictions coincide with clear phenological cues, while lower-confidence outputs correctly signal uncertainty, demonstrating reliable probability calibration.
By combining deep neural networks with agronomic expertise, our approach not only enhances classification accuracy but also empowers the model to recognize and flag its own uncertainty, an essential feature in agricultural decision‐making, where misplaced confidence can have costly consequences.
@@ 7 Conclusion
In this paper, we introduced Thermal Time‐based Temporal
Sampling (T3S), a simple, model-agnostic method that uses cumulative growing degree days to select relevant observations within satellite time series for crop growth-related applications. Based on basic ecophysiological principles of crop growth, it reduces the computational load by removing redundant information and allows to prioritize relevant and undisturbed imagery. Although we evaluated T3S with attention-based deep learning architecture, it can be seamlessly integrated with any machine learning model. When compared to other state-of-the-art approaches, T3S not only enhances predictive accuracy but also markedly improves uncertainty estimation, an aspect often overlooked in the literature. Furthermore, we made evaluation on SwissCrop, a novel, large-scale, multi-year dataset spanning three growing seasons (2021–2023) across Switzerland’s agricultural landscapes, thereby providing a real-world benchmark for compared methods. By releasing both SwissCrop and our T3S implementation to the public, we establish a robust, reproducible benchmark for the community, facilitating fair comparisons and driving further innovation in crop‐growth modelling. For future work, incorporating additional environmental drivers such as precipitation could further strengthen temporal alignment and robustness across diverse agro-climatic regions.
8 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
9 Acknowledgments
Mehmet Ozgur Turkoglu acknowledges support from Agroscope through research funding for the CropmAIpper project. Sélène Ledain acknowledges funding from the Swiss Federal Office of Agriculture and Agroscope within the Monitoring of the Swiss Agri-Environmental System (MAUS) program. This project was conducted in the context of the Swiss Agricultural Landscape Intelligence platform, funded through an Agroscope Research grant awarded to Helge Aasen.

### v3

@@ Discussion.
Figure 8: Low data regime performance showing (a) accuracy and (b) expected calibration error (ECE) averaged over six folds when training on 10% of the data, comparing T3S (Ours) and U-TAE.
@@
Figure 9: Qualitative comparison between T3S and U-TAE. Green pixels denote correct predictions, red pixels denote errors, and dark blue indicates low uncertainty, while yellow indicates high uncertainty. For best visibility, view the figure zoomed on-screen.
@@ 6 Conclusion
We present Thermal Time–based Temporal Sampling (T3S), a model-agnostic approach that leverages cumulative growing degree days to subsample satellite time series in biologically meaningful ways. By injecting this thermal-time inductive bias, T3S steers the attention toward biologically relevant periods, reducing redundancy and noise and preventing overconfident predictions under weak or contradictory evidence. We evaluated T3S on SwissCrop, a large, multi-year (2021–2023) dataset covering all of Switzerland, and release both the data and our implementation as a public benchmark. Compared to state-of-the-art methods, T3S improves both classification accuracy across years and uncertainty estimation, an aspect often overlooked in the literature. Since temperature is a main driver for many biological processes, we anticipate that T3S also has advantages for many other applications even beyond plant science. And since T3S is model-agnostic, it is applicable to any learning model. We leave these explorations for future work.
Appendix
Appendix A Training Details
The model was trained for 100 epochs using a batch size of 8 and a fixed temporal length ($T$) of 24. The input image size ($H$ and $W$) is set to $128\times 128$ pixels. We employed the Adam optimizer with an initial learning rate of 0.001, scheduled by a OneCycleLR policy with a maximum learning rate of 0.01 and a warm-up phase covering 30% of the total training steps. The U-TAE architecture is configured with an encoder having channel widths $[128,128,128,256]$, a symmetric decoder with widths $[32,64,128,256]$ and strided convolutions with kernel size $k=4$, stride $s=2$, and padding $p=1$. The temporal attention module employs $n_{\text{head}}=16$ heads and a model dimension of $d_{\text{model}}=256$. Input features are standardized to zero mean and unit variance on a per-channel basis. We implemented the model in PyTorch and trained it on a single NVIDIA RTX 4090 GPU over roughly three days. The full codebase and pretrained checkpoints will be publicly available on GitHub.
Appendix B Dataset
In this work, we work with SwissCrop 2021, 2022, 2023, a three-year crop dataset covering the entirety of Switzerland (see Figure 10) from 2021 to 2023. SwissCrop combines Sentinel-2 Level-2A bottom-of-atmosphere multi-spectral imagery, averaging 110 timestamps per season, with annual crop‐type labels for each field. The dataset includes 50 distinct crop types, each representing the primary crop grown in a field during a given season. These labels where provided by farmers to the cantons of Switzerland according to a data model provided by the Federal Office for Agriculture. [^1] https://www.blw.admin.ch/de/landwirtschaftliche-kulturflaechen Unlike many existing datasets, e.g. Pastis dataset [14], SwissCrop faithfully reflects the pronounced class imbalance found in real‐world farming: a long‐tail distribution in which a handful of major crops dominate while the majority occur only sporadically (see Figure 11).
[…]
To complement these remote sensing observations, we incorporate high-resolution gridded climate data from MeteoSwiss, which supplies near-surface temperature fields (minimum, maximum, and average) on 1 km, 2 km, and 5 km grids [11]. In our study, we use the 1 km daily minimum (TminD) and maximum (TmaxD) temperature series to compute growing degree days (GDD; refer to Section 3.3).
Appendix C Extended Related Work
Deep Learning for Crop Mapping from Satellite Imagery.
Deep learning has dramatically improved crop‐type classification from satellite time series, outperforming traditional methods based on handcrafted features [50]. Early sequence‐modeling approaches used recurrent architectures to capture phenological dynamics. Long short‐term memory (LSTM) networks demonstrated strong performance on multi‐temporal inputs [38], and the convolutional‐RNN (convRNN) version further boosted both accuracy and efficiency [51].
Attention mechanisms have emerged as a powerful alternative for modeling long‐range temporal dependencies. [41] and [15] were the first to apply self‐attention to crop classification from Sentinel-2 data and showed improved results. [44] then proposed the lightweight Temporal Attention Encoder (L-TAE), and [14] embedded L-TAE within a U-Net backbone to adaptively weigh temporal features across space and time. More recently, fully attention-based Vision Transformer (ViT) adaptations have been explored for crop mapping [2], and [33] leveraged the recently proposed Mamba state‐space architecture to improve spatio‐temporal classification performance.
To incorporate domain knowledge in deep learning architectures, several works introduce structured priors and hierarchical labels. [55] exploited multi‐scale crop class hierarchies within a parameter‐efficient convRNN (i.e, convSTAR [52]) framework, and prototype‐based supervision has been used to integrate class relationships into metric learning [45]. Positional encodings tailored to agricultural cycles have also been proposed: [31] introduced thermal-time encoding, using cumulative growing degree days, to normalize phenological shifts across regions. This is the most relevant research for this paper. Moreover, practical deployment demands handling early-season data and irregular revisit intervals. [43] added an early classification reward loss to LSTM training, improving accuracy in the season when only partial time series are available. [29] leveraged neural ordinary differential equations (NODE) to process noisy and irregularly sampled time series, and [7]’s collect–update–distribute design further addresses sampling irregularity.
Multi-Year Crop Mapping from Satellite Imagery.
Multi-year crop mapping faces challenges due to temporal domain shifts such as interannual variability in weather, phenological differences, and sensor calibration issues. These factors often degrade the performance of models trained on single-season data when applied across multiple years. [54] extended the experiments of [51] to a multi‐year setting, training and testing on data from two different years, and reporting preliminary results. [5] employed a Random Forest classifier using dense Sentinel-2 and Landsat-8 time series, monthly Sentinel-1 composites, and environmental variables to evaluate the relative contributions of optical, radar, and environmental data across three meteorologically distinct years (2017–2019). [9] demonstrated that reconstructing multi-year Sentinel-2 NDVI time series with Fourier-based fitting and applying Gradient Boosting enables improved crop mapping across different years. [13] investigated ways to optimize the use of multi-year samples in a within-season crop classification model. [32] presented a Google Earth Engine workflow that fuses field-sample data with multi-year Sentinel-2 time series to map cropping systems on smallholder farms, achieving consistently robust results across diverse landscapes. [26] benchmarked deep learning segmentation models (e.g., U-Net and SegFormer [60]) for crop mapping across multiple years, while [47] evaluated convLSTM and convSTAR [52] networks in a similar multi-year context. Very recently, [22] shows improved results for cross-year crop mapping combining a U-Net with photometric augmentation, Tversky-Focal loss, and Monte Carlo (MC) Dropout.
Uncertainty Modeling in Deep Learning.
Predictive uncertainty is commonly split into aleatoric uncertainty, which stems from inherent data variability such as class overlap or sensor noise, and epistemic uncertainty, which reflects model weight uncertainty due to insufficient coverage of the input space [21]. Exact Bayesian inference in deep networks is generally intractable, leading to approximate methods like variational inference (e.g., Bayes by Backprop [6] and Markov Chain Monte Carlo techniques such as Stochastic Gradient Langevin Dynamics [59]). However, these approaches often struggle to scale to large architectures [18]. Probabilistic ensembles, particularly deep ensembles, currently represent the gold standard for epistemic uncertainty estimation in deep networks, providing well‐calibrated and robust predictions by aggregating outputs from multiple independently trained models [23]. However, the need to train and store several large networks incurs substantial computational and memory overhead. To address this, implicit ensemble techniques, such as Monte Carlo (MC) Dropout [12], Snapshot Ensembles [20], Masksembles [10], and FiLM-Ensemble [53], approximate ensemble diversity via stochastic weight perturbations or dynamic weight combinations, significantly reducing resource requirements. Despite these gains in efficiency, implicit methods generally underperform full ensembles in both predictive accuracy and uncertainty calibration. More recently, LoRA-Ensemble has been introduced as a parameter‐efficient alternative for transformer‐based architectures, achieving uncertainty estimates competitive with deep ensembles at a fraction of the cost [19].
Uncertainty Modeling in Satellite Imagery and Crop Mapping
Uncertainty quantification has become a standard component of satellite-based land-cover and crop-mapping workflows, as it provides essential information for risk-aware decision-making in agricultural management and ecosystem monitoring. Generating per-pixel or per-field uncertainty maps alongside deterministic predictions enables practitioners to identify low-confidence areas and tailor downstream analyses accordingly. For example, [24] employs a deep ensemble of convolutional neural networks to estimate global canopy height from sparse LiDAR samples and multi-sensor satellite imagery, producing uncertainty intervals that reflect both model and data variability. [4] extends this ensemble approach to country-scale forest structure estimation at 10m resolution, demonstrating that dense uncertainty fields help flag regions where auxiliary field data are most needed.
In the context of crop yield and type mapping, several recent studies have adopted stochastic inference techniques. [49] introduces an attention-guided multi-level crop network for winter wheat yield estimation at the county scale, applying MC Dropout during inference to derive spatially explicit uncertainty maps of yield forecasts. Similarly, [22] leverages MC Dropout in a U-Net segmentation framework to quantify per-pixel uncertainty in annual cropland masks across multiple years in Ghana, using the resulting confidence estimates to discard unreliable predictions. [27] integrates MC Dropout in a ViT–ResNet feature encoder to produce field-level uncertainty scores, filtering out low-confidence labels to improve overall mapping precision. Beyond dropout-based methods, [3] proposes a probabilistic crop-type mapping method in which an ensemble of classifiers, each trained with perturbed inputs and hyperparameters on Sentinel-1, Sentinel-2, and Landsat-8 data, generates multiple crop maps over Germany; per-pixel frequency counts across the ensemble yield robust probability distributions and uncertainty metrics. [28] refines the 30m cropland data by harnessing its built-in Random Forest confidence layer: low-certainty regions are iteratively reclassified using targeted ancillary datasets, resulting in a more accurate and self-aware mapping product. Collectively, these approaches demonstrate the critical role of uncertainty modeling in improving the reliability and interpretability of satellite-driven crop and canopy mapping systems.